OpenStudy (anonymous):

solve \[\left(\begin{matrix}x_1 \\ x_2\\x_3\\x_4\end{matrix}\right)'=\left[\begin{matrix}0 & -1 & 0 & 0\\ -2 & -2 & 1 & 0\\ 0 & 0 & 0& 1\\1&0&-2&-2\end{matrix}\right]\left(\begin{matrix}x_1 \\ x_2\\x_3\\x_4\end{matrix}\right)\]

OpenStudy (anonymous):

@satellite73 can you help?

OpenStudy (anonymous):

@terenzreignz can you help?

terenzreignz (terenzreignz):

That prime symbol means transpose?

OpenStudy (anonymous):

no it means its derivative... it's an ODE problem

terenzreignz (terenzreignz):

Okay... never seen these before D: mercy T.T I'll still look into it though, try to enlighten myself ^_^

OpenStudy (anonymous):

ok thanks. i found the eigenvalues as such:\[\lambda=(-1+\sqrt{2}i),(-1-\sqrt{2}i),(-1),(-1)\]I don't know how to find the eigenvectors

terenzreignz (terenzreignz):

I'm allergic to matrices -_- But finding eigenvectors for a specific eigenvalue seems simple enough, in the case where the eigenvalue is -1, just look for the vectors which satisfy \[\left[\begin{matrix}0 & -1 & 0 & 0\\ -2 & -2 & 1 & 0\\ 0 & 0 & 0& 1\\1&0&-2&-2\end{matrix}\right]\left(\begin{matrix}x_1 \\ x_2\\x_3\\x_4\end{matrix}\right)=-\left(\begin{matrix}x_1 \\ x_2\\x_3\\x_4\end{matrix}\right)\]

OpenStudy (anonymous):

i don't have a clue how to do the 4x4 at all :(